# Not All Latent Spaces Are Flat: Hyperbolic Concept Control

### 14 Mar 2026 v1

@@ 1 Introduction
The rapid progress of generative models, particularly diffusion models, has enabled high-quality image synthesis from natural language prompts, substantially lowering the barrier to visual content creation. Beyond prompt engineering, a growing line of work has explored concept control by steering, i.e., post-hoc manipulation of model representations to encourage or suppress specific semantic attributes during generation (Dathathri et al., 2020; Lee et al., 2025; Schramowski et al., 2023; Yoon et al., 2025). Most existing steering approaches operate in Euclidean embedding spaces inherited from vision–language models (VLMs) such as CLIP (Radford et al., 2021), and rely on linear manipulations of text or latent representations, including inference-time vector steering (Schramowski et al., 2023; Yoon et al., 2025), direct weight editing (Gandikota et al., 2023; Li et al., 2024), or activation-level modulation (Zhang et al., 2024). While these methods provide flexible control without retraining, they offer limited structural guarantees on how semantic changes evolve, making their behavior difficult to predict and regulate under substantial interventions.
@@
In this work, we argue that these limitations stem from the geometry of the representation space itself. Semantic concepts exhibit hierarchical relationships, where different levels of abstraction and semantic inclusion structure the representation space. An embedding space that explicitly reflects such structure can support smoother transitions, in which semantic changes are introduced progressively while remaining within coherent semantic regions. Hyperbolic geometry provides a natural substrate for such representations, as it enables hierarchical organization through its geometric structure (Nickel and Kiela, 2017; Ganea et al., 2018b; Vilnis et al., 2018). Recent hyperbolic VLMs, such as MERU (Desai et al., 2023) and HyCoCLIP (Pal et al., 2025), demonstrate that images and text can be embedded in spaces where semantic entailment and specificity are explicitly encoded.
Building on these insights, we introduce HyCon, a hyperbolic concept control framework for text-to-image generative models. Our approach operates in the hyperbolic embedding space learned by HyCoCLIP (Pal et al., 2025), where concepts are organized via entailment relations and represented as entailment cones. Steering in this space corresponds to traversing the embedding geometry, as shown in Figure 1, in a way that respects the underlying hierarchical concept structure, rather than applying arbitrary linear offsets. To integrate hyperbolic control with existing generative models, we employ a lightweight logarithmic adapter that maps hyperbolic text embeddings to the conditioning spaces of pretrained diffusion backbones, without retraining the generative model. We demonstrate that hyperbolic control by steering enables more stable and predictable semantic manipulation compared to Euclidean baselines, particularly as the control strength varies. The effect of steering in HyCon follows the hierarchical organization encoded in the embedding geometry, rather than relying on heuristic scaling of linear directions. We evaluate our method across multiple modern diffusion backbones, including Stable Diffusion 3 and 3.5 (Esser et al., 2024), SDXL (Podell et al., 2023), and FLUX (BlackForestLabs, 2024), and validate its effectiveness in both retrieval and generative settings. Our contributions are summarized as follows:
@@ 2 Related work
Control in Text-to-Image Generation. Controlling the behavior of diffusion models (DMs) has always been a central problem in conditional image generation. Early approaches relied on classifier guidance, where gradients from an external classifier are used to steer the denoising process toward desired classes or away from undesired content (Dhariwal and Nichol, 2021). This idea was later refined by classifier-free guidance, which enables a controllable trade-off between fidelity and diversity without requiring auxiliary classifiers (Ho and Salimans, 2022).
@@
Beyond guidance mechanisms, recent methods explored control by directly manipulating conditioning signals or internal representations. Prompt-to-Prompt (Hertz et al., 2022) enables localized image edits by modifying cross-attention maps associated with specific tokens, allowing text-driven edits without retraining. Textual Inversion (Gal et al., 2022) and DreamBooth (Ruiz et al., 2023) personalize generation by learning new text embeddings or finetuning the model for specific concepts. While effective, these approaches typically require per-concept optimization and do not provide continuous, interpretable control at inference time. A seminal work added a few parameters to control by adding spatial constraints using ControlNet (Zhang et al., 2023a).
@@
Latent Steering via Concept Vectors. A growing body of research focuses on controllability by directly manipulating latent or intermediate representations of DMs. Several approaches identify semantic directions in the latent spaces, corresponding to undesirable or sensitive concepts, and use them to control generation at inference. SLD (Schramowski et al., 2023) suppresses unsafe concepts during the denoising, while SAFREE (Yoon et al., 2025) explicitly constructs an unsafe subspace in the text embedding space, projecting prompts away from it. Other methods pursue related objectives using different mechanisms. UCE (Gandikota et al., 2023) performs closed-form weight editing to erase concepts from DMs, while RECE (Li et al., 2024) combines adversarial finetuning with analytical corrections to improve robustness. ANT (Zhang et al., 2024) dynamically steers the sampling trajectory away from undesired regions. These approaches demonstrate that Euclidean latent spaces encode semantically meaningful directions that can be exploited for control and safety. However, such directions lack an explicit notion of hierarchy or graded semantic inclusion, limiting interpretability and stability.
@@
Hyperbolic Representations. Hyperbolic geometry has been widely studied as a representation space for hierarchical and taxonomic data. Foundational work on Poincaré embeddings (Nickel and Kiela, 2017) and entailment cones (Ganea et al., 2018a) shows that hyperbolic spaces naturally encode partial orders and concept specificity, offering lower distortion than Euclidean embeddings when representing tree-like structures, such as hierarchies. These ideas have recently been extended to hyperbolic VLMs such as MERU (Desai et al., 2023), ATMG (Ramasinghe et al., 2024), and LVH (Wang et al., 2024). The recent HyCoCLIP (Pal et al., 2025) embeds images and text into a shared hyperbolic space structured by entailment relations, where more general concepts lie closer to the origin, while more specific ones are positioned toward the boundary. Although hyperbolic VLMs have initially been studied for retrieval and compositional understanding, their geometric structure suggests new opportunities for semantic control, which we explore in this work. We leverage hyperbolic representations not as a replacement for existing generative models, but as a control layer for structured manipulation of prompt embeddings. By operating in a space where semantic specificity and inclusion are explicitly encoded, we use hyperbolic representations as a principled substrate for defining and applying steering operations.
@@ 3 Motivation
Are entailment cones semantically consistent? Semantic concepts are inherently compositional, often sharing common properties or attributes. Hyperbolic geometry provides a natural space for capturing these relationships through the structure of entailment cones (Ganea et al., 2018b). Indeed, in models like HyCoCLIP (Pal et al., 2025), concepts are represented as conical regions whose volumes and their positions relative to the origin reflect the specificity of the concept. Crucially, the intersection of these cones defines a semantically meaningful overlap, representing the common ground between distinct categories.
@@ 4.1 Preliminaries
Refusal Vectors. Euclidean latent steering typically achieves semantic control by editing embeddings along concept-specific directions (Mikolov et al., 2013; Marks and Tegmark, 2023; Park et al., 2024; Facchiano et al., 2026). This relies on the assumption that high-level concepts are (at least locally) linearly encoded in representation space, as formalized by Concept Activation Vectors (TCAV) (Kim et al., 2018). Within this view, a refusal vector is a direction $\mathbf{v}_{c}\in\mathbb{R}^{d}$ associated with an undesired concept $c$ such that moving an embedding orthogonally to $\mathbf{v}$ reduces that concept in the generated output (Arditi et al., 2025). Given a new input with embedding $\mathbf{x}\in\mathbb{R}^{d}$ and a refusal vector $\mathbf{v}_{c}$ steering is typically performed as:
@@ 5.3 Steering Experiments
Setup. We evaluate hyperbolic concept control in generative settings by interfacing HyCon with multiple pretrained backbones, including SDXL (Podell et al., 2023), SD 3, SD 3.5 (Esser et al., 2024), and FLUX (BlackForestLabs, 2024; BlackForestLabs et al., 2025). We benchmark on four established safety-oriented datasets: P4D (Chin et al., 2024), Ring-A-Bell (Tsai et al., 2024), MMA-Diffusion (Yang et al., 2024), and UnlearnDiffAttk (Zhang et al., 2023b). Image quality is additionally evaluated on a subset of MS-COCO prompts. We compare hyperbolic control against SAFREE (Yoon et al., 2025), which removes unsafe concepts via orthogonal projection in the Euclidean space. SAFREE is reported only for DMs that expose compatible text-conditioning interfaces (thus, FLUX is omitted). Following prior art (Yoon et al., 2025), we report safety-related metrics, including NudeNet (notAI tech, 2019) and GPT-based classifiers where applicable. Image quality and fidelity on the COCO retain set are measured using FID, CLIPScore, and LPIPS.

### 03 Apr 2026 v3

@@ 1 Introduction
The rapid progress of generative models, particularly diffusion models, has enabled high-quality image synthesis from natural language prompts, substantially lowering the barrier to visual content creation. Beyond prompt engineering, a growing line of work has explored concept control by steering, i.e., post-hoc manipulation of model representations to encourage or suppress specific semantic attributes during generation (Dathathri et al., 2020; Lee et al., 2025; Schramowski et al., 2023; Yoon et al., 2025). Most existing steering approaches operate in Euclidean embedding spaces inherited from vision–language models (VLMs) such as CLIP (Radford et al., 2021), and rely on linear manipulations of text or latent representations, including inference-time vector steering (Schramowski et al., 2023; Yoon et al., 2025), direct weight editing (Gandikota et al., 2023; Gong et al., 2024), or activation-level modulation (Li et al., 2025). While these methods provide flexible control without retraining, they offer limited structural guarantees on how semantic changes evolve, making their behavior difficult to predict and regulate under substantial interventions.
@@
In this work, we argue that these limitations stem from the geometry of the representation space itself. Semantic concepts exhibit hierarchical relationships, where different levels of abstraction and semantic inclusion structure the representation space. An embedding space that explicitly reflects such structure can support smoother transitions, in which semantic changes are introduced progressively while remaining within coherent semantic regions. Hyperbolic geometry provides a natural substrate for such representations, as it enables hierarchical organization through its geometric structure (Nickel and Kiela, 2017; Ganea et al., 2018a; Vilnis et al., 2018). Recent hyperbolic VLMs, such as MERU (Desai et al., 2023) and HyCoCLIP (Pal et al., 2025), demonstrate that images and text can be embedded in spaces where semantic entailment and specificity are explicitly encoded.
Building on these insights, we introduce HyCon, a hyperbolic concept control framework for text-to-image generative models. Our approach operates in the hyperbolic embedding space learned by HyCoCLIP (Pal et al., 2025), where concepts are organized via entailment relations and represented as entailment cones. Steering in this space corresponds to traversing the embedding geometry, as shown in Figure 1, in a way that respects the underlying hierarchical concept structure, rather than applying arbitrary linear offsets. To integrate hyperbolic control with existing generative models, we employ a lightweight logarithmic adapter that maps hyperbolic text embeddings to the conditioning spaces of pretrained diffusion backbones, without retraining the generative model. We demonstrate that hyperbolic control by steering enables more stable and predictable semantic manipulation compared to Euclidean baselines, particularly as the control strength varies. The effect of steering in HyCon follows the hierarchical organization encoded in the embedding geometry, rather than relying on heuristic scaling of linear directions. We evaluate our method across multiple modern diffusion backbones, including Stable Diffusion 3 and 3.5 (Esser et al., 2024), SDXL (Podell et al., 2024), and FLUX (BlackForestLabs et al., 2025), and validate its effectiveness in both retrieval and generative settings. Our contributions are summarized as follows:
@@ 2 Related work
Control in Text-to-Image Generation. Controlling the behavior of diffusion models (DMs) has always been a central problem in conditional image generation. Early approaches relied on classifier guidance, where gradients from an external classifier are used to steer the denoising process toward desired classes or away from undesired content (Dhariwal and Nichol, 2021). This idea was later refined by classifier-free guidance, which enables a controllable trade-off between fidelity and diversity without requiring auxiliary classifiers (Ho and Salimans, 2021).
@@
Beyond guidance mechanisms, recent methods explored control by directly manipulating conditioning signals or internal representations. Prompt-to-Prompt (Hertz et al., 2023) enables localized image edits by modifying cross-attention maps associated with specific tokens, allowing text-driven edits without retraining. Textual Inversion (Gal et al., 2023) and DreamBooth (Ruiz et al., 2023) personalize generation by learning new text embeddings or finetuning the model for specific concepts. While effective, these approaches typically require per-concept optimization and do not provide continuous, interpretable control at inference time. A seminal work added a few parameters to control by adding spatial constraints using ControlNet (Zhang et al., 2023).
@@
Latent Steering via Concept Vectors. A growing body of research focuses on controllability by directly manipulating latent or intermediate representations of DMs. Several approaches identify semantic directions in the latent spaces, corresponding to undesirable or sensitive concepts, and use them to control generation at inference. SLD (Schramowski et al., 2023) suppresses unsafe concepts during the denoising, while SAFREE (Yoon et al., 2025) explicitly constructs an unsafe subspace in the text embedding space, projecting prompts away from it. Other methods pursue related objectives using different mechanisms. UCE (Gandikota et al., 2023) performs closed-form weight editing to erase concepts from DMs, while RECE (Gong et al., 2024) combines adversarial finetuning with analytical corrections to improve robustness. ANT (Li et al., 2025) dynamically steers the sampling trajectory away from undesired regions. These approaches demonstrate that Euclidean latent spaces encode semantically meaningful directions that can be exploited for control and safety. However, such directions lack an explicit notion of hierarchy or graded semantic inclusion, limiting interpretability and stability.
@@
Hyperbolic Representations. Hyperbolic geometry has been widely studied as a representation space for hierarchical and taxonomic data. Foundational work on Poincaré embeddings (Nickel and Kiela, 2017) and entailment cones (Ganea et al., 2018b) shows that hyperbolic spaces naturally encode partial orders and concept specificity, offering lower distortion than Euclidean embeddings when representing tree-like structures, such as hierarchies. These ideas have recently been extended to hyperbolic VLMs such as MERU (Desai et al., 2023), ATMG (Ramasinghe et al., 2024), and LVH (Wang et al., 2025). The recent HyCoCLIP (Pal et al., 2025) embeds images and text into a shared hyperbolic space structured by entailment relations, where more general concepts lie closer to the origin, while more specific ones are positioned toward the boundary. Although hyperbolic VLMs have initially been studied for retrieval and compositional understanding, their geometric structure suggests new opportunities for semantic control, which we explore in this work. We leverage hyperbolic representations not as a replacement for existing generative models, but as a control layer for structured manipulation of prompt embeddings. By operating in a space where semantic specificity and inclusion are explicitly encoded, we use hyperbolic representations as a principled substrate for defining and applying steering operations.
@@ 3 Motivation
Are entailment cones semantically consistent? Semantic concepts are inherently compositional, often sharing common properties or attributes. Hyperbolic geometry provides a natural space for capturing these relationships through the structure of entailment cones (Ganea et al., 2018a). Indeed, in models like HyCoCLIP (Pal et al., 2025), concepts are represented as conical regions whose volumes and their positions relative to the origin reflect the specificity of the concept. Crucially, the intersection of these cones defines a semantically meaningful overlap, representing the common ground between distinct categories.
@@ 4.1 Preliminaries
Refusal Vectors. Euclidean latent steering typically achieves semantic control by editing embeddings along concept-specific directions (Mikolov et al., 2013; Marks and Tegmark, 2024; Park et al., 2024; Facchiano et al., 2026). This relies on the assumption that high-level concepts are (at least locally) linearly encoded in representation space, as formalized by Concept Activation Vectors (TCAV) (Kim et al., 2018). Within this view, a refusal vector is a direction $\mathbf{v}_{c}\in\mathbb{R}^{d}$ associated with an undesired concept $c$ such that moving an embedding orthogonally to $\mathbf{v}$ reduces that concept in the generated output (Arditi et al., 2025). Given a new input with embedding $\mathbf{x}\in\mathbb{R}^{d}$ and a refusal vector $\mathbf{v}_{c}$ steering is typically performed as:
@@ 5.3 Steering Experiments
Setup. We evaluate hyperbolic concept control in generative settings by interfacing HyCon with multiple pretrained backbones, including SDXL (Podell et al., 2024), SD 3, SD 3.5 (Esser et al., 2024), and FLUX (BlackForestLabs et al., 2025). We benchmark on four established safety-oriented datasets: P4D (Chin et al., 2024), Ring-A-Bell (Tsai et al., 2023), MMA-Diffusion (Yang et al., 2024), and UnlearnDiffAttk (Zhang et al., 2024). Image quality is additionally evaluated on a subset of MS-COCO prompts. We compare hyperbolic control against SAFREE (Yoon et al., 2025), which removes unsafe concepts via orthogonal projection in the Euclidean space. SAFREE is reported only for DMs that expose compatible text-conditioning interfaces (thus, FLUX is omitted). Following prior art (Yoon et al., 2025), we report safety-related metrics, including NudeNet (notAI tech, 2019) and GPT-based classifiers where applicable. Image quality and fidelity on the COCO retain set are measured using FID, CLIPScore, and LPIPS.